OpenStudy (anonymous):

Find the equation of the parabola with focus (5, 1) and directrix y = -1.

OpenStudy (anonymous):

A) (x - 5)2 = 4y B) (y - 5)2 = 4x C) (x - 5)2 = -4y D) (y - 5)2 = -4x

OpenStudy (anonymous):

the 2's are supposed to be squared